# First-passage properties of the jump process with a drift. The general case

### v1

@@ 7. Moments of $\tau$ and $n$
Having derived rather cumbersome expressions for the asymptotic coefficients, it is imperative to test their validness. The first check is to compare them against the exactly solvable case where both $p(t)$ and $q(M)$ are exponentially distributed, for which exact analytical expressions were obtained in [36]. A motivated reader can verify that the asymptotic expressions do coincide with these exact results.
Figure 11. Comparison between the numerical simulations and the asymptotic expansions of the mean values of $\tau$ and $n$ for the distributions (187). The parameters of the model are $\alpha=2$, $\gamma=1$, $\ell=0{.}9$. Black dashed lines correspond to asymptotic expansions for $X_{0}\to\infty$. Blue dash-dotted lines correspond to asymptotic expansions for $X_{0}\to 0$. Red solid lines correspond to the numerical simulation, which was obtained be generating $10^{8}$ trajectories for 5050 values of $X_{0}$ in $[0;2]$
However, testing the asymptotics only against this special case would not be prudent. In the double exponential case, the mean and variance are actually linear functions of $X_{0}$, which means that the asymptotic behaviors for $X_{0}\to\infty$ and $X_{0}\to 0$ are essentially identical. More importantly, this might mislead the reader into believing that such linear behavior is generic. This suggests that we need to test the asymptotic expansions against a more general case of the distributions $p(t)$ and $q(M)$, making numerical simulations the natural choice. Following this idea, we performed Monte Carlo simulations for a specific process where the time intervals are drawn from a half-Gaussian distribution and the jumps follow a uniform distribution:
$q(M)=\sqrt{\frac{2}{\pi}}\gamma\;e^{-\frac{1}{2}\gamma^{2}M^{2}},\qquad p(t)=\begin{cases}\frac{1}{\ell},&0<t\leq\ell,\\ 0,&t>\ell.\end{cases}$ (187)
@@
$\left\langle M-\alpha t\right\rangle=\frac{1}{\gamma}\sqrt{\frac{2}{\pi}}-\frac{\alpha\ell}{2}.$ (188)
The specific choice (187) illustrates two key features: the support of $p(t)$ is bounded rather than infinite, and this combination appears to have no analytical solution — neither the auxiliary integrals nor the corresponding renewal equations have known closed forms.
@@ 8. Conclusion
Our main result is confirming that the intuitive picture obtained in [36] for exactly solvable cases extends to the general setting. We identified three regimes: the survival regime (weak drift) where the process has finite probability of remaining positive indefinitely, the absorption regime (strong drift) where first-passage occurs with certainty, and the critical point separating these regimes. In both survival and absorption regimes, survival probabilities approach their asymptotic values exponentially fast, with explicitly computable decay rates. At the critical point, exponential decay is replaced by algebraic decay with universal scaling behavior described by error functions.
@@
A natural question concerns the case where our assumptions are violated. As mentioned earlier, the smooth density requirement is largely technical and can likely be relaxed without substantial modifications to the analysis. The case of heavy-tailed distributions, where the second moment of the effective random walk is infinite, on the other hand, presents a challenge. Allowing the distributions to be heavy-tailed would alter the behavior of $F(k;\rho)$ in ways that could invalidate our asymptotic analysis. The heavy-tailed case therefore merits detailed investigation. Another important assumption is that $t_{i}$ and $M_{j}$ are independent. It would be interesting to study whether the proposed formalism can be generalized to some extent to situations where correlations are present, as was done for the Sparre-Andersen theorem in [64].
@@ Appendix A Asymptotic expansions for the means and variances
This appendix is devoted to the computation of the asymptotic expansions of the means and variances of $n$ and $\tau$. First we extract the asymptotic behavior of $I(\lambda;\rho,s)$ and its two first derivatives using the approach of [63]. A key ingredient is the expression (164) that allows us to construct the asymptotic expansion for the Mellin convolution integrals (162) from the asymptotic behavior of the convoluted functions (163). These results are then used to construct the expansions of the mean and variances given by (168), (169), and (170).
@@ A.2. Expansion of $\partial_{s}I(\lambda;0,s)$
The transform of $g_{1}$ is however more subtle, as the integral (165) diverges, and hence the Mellin transform should be understood in the analytical continuation sense. The procedure is explained in [63]; we state only the result here
@@ A.3. Expansion of $\partial_{\rho}I(\lambda;\rho,1)$
$a_{0}=\frac{1}{\alpha}\frac{\left\langle M(M-\alpha t)^{2}\right\rangle}{\left\langle(M-\alpha t)^{2}\right\rangle},\quad\alpha_{0}=0,\quad\alpha_{1}=1,b_{0}=1,\quad\beta_{0}=0,\quad\beta_{1}=2$ (271)
@@
As for the Mellin transform of $r_{1}$, we face the same problem as for $g_{1}$ in (230). Specifically, the integral (165) diverges and we need to perform the analytic continuation as prescribed in [63]. The resulting expression is
@@ A.4. Asymptotic expansion for the moments
$\left.\frac{\partial^{2}}{\partial\rho^{2}}\hat{Q}(\rho,1\,|\,\lambda)\right|_{\rho=0}=\frac{2}{\alpha^{2}\lambda^{3}}\left(1-\frac{\langle M\rangle}{\langle M-\alpha t\rangle}\left[1-\frac{\langle M\rangle}{\langle M-\alpha t\rangle}\right]e^{I(\lambda;0,1)}\right)\\ +\frac{2\,\langle M\rangle}{\alpha^{2}\lambda^{2}}\left(\frac{2\langle M(M-\alpha t)\rangle}{\langle M-\alpha t\rangle^{2}}-\langle M\rangle\frac{\langle(M-\alpha t)^{2}\rangle}{\langle M-\alpha t\rangle^{3}}\right)e^{I(\lambda;0,1)}\\ +\frac{2}{\alpha^{2}\lambda^{2}}\frac{\langle M\rangle}{\langle M-\alpha t\rangle}\left(\left.\frac{\partial}{\partial\rho}I(\lambda;\rho,1)\right|_{\rho=0}-\frac{\langle M^{2}\rangle}{\langle M\rangle}\right)e^{I(\lambda;0,1)}.$ (301)

### v2

@@ 8. Moments of $\tau$ and $n$
Having derived rather cumbersome expressions for the asymptotic coefficients, we must test their validity. The first check is to compare them against the exactly solvable case where both $p(t)$ and $q(M)$ are exponentially distributed, for which the analytical expressions were obtained in [undefal]. A motivated reader can verify that the asymptotic expressions do coincide with these exact results.
Figure 17. Comparison between direct Monte Carlo simulations (black circles, $10^{8}$ independent trajectory realizations for each value of $X_{0}$) and the asymptotic expansions (189),(190), (191), and (192) for the means and the variances of $\tau$ and $n$. The simulations use distributions (209): uniform inter-arrival times and half-Gaussian jump amplitudes. Parameters: $\alpha=2$, $\gamma=1$, $\ell=0{.}9$. Red dashed lines show the large-$X_{0}$ expansions; blue dotted lines show the small-$X_{0}$ expansions.
However, testing the asymptotics only against this special case would not be prudent. In the double exponential case, the mean and variance are linear functions of $X_{0}$, which means that the asymptotic behaviors for $X_{0}\to\infty$ and $X_{0}\to 0$ are essentially identical. More importantly, this might mislead the reader into believing that such linear behavior is generic. This suggests that we need to test the asymptotic expansions against a more general case of the distributions $p(t)$ and $q(M)$, making numerical simulations the natural choice.
Unlike the previous sections, where we chose inverse-Gaussian distributions for their explicit decay rates, here the auxiliary integrals must be evaluated numerically in any case. We therefore take this opportunity to test against a different family of distributions, further demonstrating the generality of the approach. Specifically, we perform Monte Carlo simulations for a process where the time intervals are drawn from a half-Gaussian distribution and the jumps follow a uniform distribution:
$p(t)=\begin{cases}\frac{1}{\ell},&0<t\leq\ell,\\ 0,&t>\ell,\end{cases}\qquad q(M)=\sqrt{\frac{2}{\pi}}\gamma\;e^{-\frac{1}{2}\gamma^{2}M^{2}},$ (209)
@@
$\left\langle M-\alpha t\right\rangle=\frac{1}{\gamma}\sqrt{\frac{2}{\pi}}-\frac{\alpha\ell}{2},\qquad\alpha_{c}=\frac{2}{\gamma\ell}\sqrt{\frac{2}{\pi}}.$ (210)
The choice (209) illustrates two key features: (i) the support of $p(t)$ is bounded rather than infinite; (ii) this combination appears to have no analytical solution — neither the auxiliary integrals nor the corresponding renewal equations have known closed forms.
@@ 9. Conclusion
Our main result is confirming that the intuitive picture obtained in [undefal] for exactly solvable cases extends to the general setting. We identified three regimes: the survival regime (weak drift) where the process has finite probability of remaining positive indefinitely, the absorption regime (strong drift) where first-passage occurs with certainty, and the critical point separating these regimes. In both survival and absorption regimes, survival probabilities approach their asymptotic values exponentially fast, with explicitly computable decay rates. At the critical point, exponential decay is replaced by algebraic decay with universal scaling behavior described by error functions.
@@
A natural question concerns the case where our assumptions are violated. As mentioned earlier, the smooth density requirement is largely technical and can likely be relaxed without substantial modifications to the analysis. The case of heavy-tailed distributions, where the second moment of the effective random walk is infinite, on the other hand, presents a challenge. Allowing the distributions to be heavy-tailed would alter the behavior of $F(k;\rho)$ in ways that could invalidate our asymptotic analysis. The heavy-tailed case therefore merits detailed investigation. Another important assumption is that $t_{i}$ and $M_{j}$ are independent. It would be interesting to study whether the proposed formalism can be generalized to some extent to situations where correlations are present, as was done for the Sparre-Andersen theorem in [undefaao].
@@ Acknowledgement
Data availability statement
No new data were created or analyzed in this study.
@@ Appendix A Asymptotic expansions for the means and variances
This appendix is devoted to the computation of the asymptotic expansions of the means and variances of $n$ and $\tau$. First we extract the asymptotic behavior of $I(\lambda;\rho,s)$ and its two first derivatives using the approach of [undefaan]. A key ingredient is the expression (185) that allows us to construct the asymptotic expansion for the Mellin convolution integrals (183) from the asymptotic behavior of the convoluted functions (184). These results are then used to construct the expansions of the mean and variances given by (189), (191), and (192).
@@ A.2. Expansion of $\partial_{s}I(\lambda;0,s)$
The transform of $g_{1}$ is however more subtle, as the integral (186) diverges, and hence the Mellin transform should be understood in the analytical continuation sense. The procedure is explained in [undefaan]; we state only the result here
@@ A.3. Expansion of $\partial_{\rho}I(\lambda;\rho,1)$
$a_{0}=\frac{1}{\alpha}\frac{\left\langle M(M-\alpha t)^{2}\right\rangle}{\left\langle(M-\alpha t)^{2}\right\rangle},\quad\alpha_{0}=0,\quad\alpha_{1}=1,\quad b_{0}=1,\quad\beta_{0}=0,\quad\beta_{1}=2$ (293)
@@
As for the Mellin transform of $r_{1}$, we face the same problem as for $g_{1}$ in (252). Specifically, the integral (186) diverges and we need to perform the analytic continuation as prescribed in [undefaan]. The resulting expression is
@@ A.4. Asymptotic expansion for the moments
$\left.\frac{\partial^{2}}{\partial\rho^{2}}\hat{Q}(\rho,1\,|\,\lambda)\right|_{\rho=0}=\frac{2}{\alpha^{2}\lambda^{3}}\left(1-\frac{\langle M\rangle}{\langle M-\alpha t\rangle}\left[1-\frac{\langle M\rangle}{\langle M-\alpha t\rangle}\right]e^{I(\lambda;0,1)}\right)\\ \frac{1}{\lambda^{2}}\frac{\langle t\rangle^{2}\big[\left\langle M^{2}\right\rangle-\langle M\rangle^{2}\big]+\langle M\rangle^{2}\big[\left\langle t^{2}\right\rangle-\left\langle t\right\rangle^{2}\big]}{\left\langle M-\alpha t\right\rangle^{3}}e^{I(\lambda;0,1)}\\ +\frac{2}{\alpha\lambda^{2}}\frac{\langle M\rangle}{\langle M-\alpha t\rangle}e^{I(\lambda;0,1)}\left.\frac{\partial}{\partial\rho}I(\lambda;\rho,1)\right|_{\rho=0}.$ (323)